# LingoQA: Video Question Answering for Autonomous Driving

### v2

@@ Abstract
We introduce LingoQA, a novel dataset and benchmark for video question answering in autonomous driving. The dataset contains 28K unique scenarios and 419K annotations. Evaluating state-of-the-art vision-language models on our benchmark shows that their performance is below human capabilities, with GPT-4V responding truthfully to 56.67% of the questions compared to 93.4% for humans. For evaluation, in addition to conducting a human study, we propose a truthfulness classifier, called Lingo-Judge, that achieves a 0.95 Spearman correlation coefficient to human evaluations, surpassing existing techniques like METEOR, BLEU, CIDEr, and GPT-4. We establish a baseline vision-language model and run extensive ablation studies to understand its performance. We release our dataset and benchmark [^1] https://github.com/wayveai/LingoQA and hope that it will provide a thorough evaluation platform for future vision-language models in autonomous driving.
{strip}
Figure 1: LingoQA is a comprehensive benchmark for Video Question Answering in autonomous driving. Our baseline vision-language model on this benchmark, can answer questions related to driving reasoning, object recognition, action justification, and scene description.
@@ 1 Introduction
Our focus is on vision-only end-to-end autonomous driving, aiming to bridge the gap between data-driven decision-making and user trust. We introduce LingoQA, a benchmark designed for autonomous driving video QA, utilizing a novel dataset comprising more than 419k QA pairs. Distinguished by its free-form approach to questions and answers, this dataset broadens the scope of autonomous driving video QA, encompassing reasoning and action justifications. Additionally, we publish a comprehensive evaluation suite consisting of 1,000 examples. At the core of our benchmark lies a novel evaluation metric based on a learned text classifier called Lingo-Judge, inspired by GPT-Judge used in TruthfulQA [34]. We perform rigorous studies correlating automatic metrics to human preferences and find that Lingo-Judge achieves a 0.950 Spearman and 0.993 Pearson correlation coefficient, surpassing existing automated labelling techniques like METEOR [5], BLEU [40], CIDEr [49], and GPT-4 [39] on our benchmark, while being fast enough for frequent runs during training and development. The evaluation code and the weights for the classifier will be released with the paper to support robust benchmarking video question answering in autonomous driving.
@@
  • LingoQA Benchmark: We introduce a novel benchmark for autonomous driving video QA using a learned text classifier for evaluation. It outperforms existing metrics, including GPT-4, with a Spearman coefficient of 0.950 indicating a strong correlation with human evaluation.
  • LingoQA Dataset: Our 419.9k QA pair dataset stands out with its free-form questions and answers, covering not just perception but also driving reasoning from the drivers directly, broadening the scope of autonomous driving video QA.
  • LingoQA Baseline: Through testing of various video-language components on LingoQA, we find that the most effective approach involves partially fine-tuning the attention layers of our vision-language model equipped with Vicuna-1.5-7B [13] and a late video fusion technique. We establish a new baseline for this field with an identified model combination. Example outputs from the model are shown in Figure LingoQA: Video Question Answering for Autonomous Driving.
@@ 2.2 Evaluation Metrics
Progress has been relatively slow in developing vision-language models for autonomous driving, with only a few works aiming to quantitatively improve upon prior work [28, 27, 56]. A key challenge consists of automated, reproducible evaluation metrics that are highly correlated with human ratings, particularly due to the inherent complexities in assessing natural language. ADAPT [27] reports human feedback in addition to standard natural language generation metrics, while DriveGPT [56] reports ChatGPT ratings. Automated methods such as BLEU [40], METEOR [5], ROUGE [33] show weak alignment with human feedback [49]. CIDEr [49] is also based on n-gram level similarity, as opposed to capturing the correctness of an answer based on its meaning. Newer evaluation metrics using ChatGPT have shown improvement in the area of sentence understanding, while still having limitations, such as providing high scores to elaborate, eloquent, but incorrect sentences [2]. Evaluation based on human feedback is subjective and difficult to reproduce. In this work, we address this challenge by introducing a novel video QA benchmark for autonomous driving that checks for factual correctness and is highly correlated to human correctness ratings on our proposed evaluation dataset.
@@ 2.3 Datasets for Autonomous Driving
Our proposed dataset LingoQA addresses the existing gap in autonomous driving as it contains a diverse set of questions related to driving behaviours and scenery in addition to perception questions related to object presence and positioning. The evaluation dataset probes areas such as description, counting, localisation, anticipation, attention, and action justification. This dataset has the strength of being diverse with respect to the language used while being grounded in human reasoning. Examples of the complex questions and answers existent in the dataset are provided in Appendix A.
@@ 3 LingoQA Benchmark
In this section, we introduce LingoQA, a benchmark to evaluate video question-answering models for autonomous driving.
3.1 Evaluation Dataset
We collected a small, low-density dataset from in-house human labelers, creating both the questions and the answers associated with the short videos. We labeled a small portion of held-out data on 500 human-generated questions using 20+ different evaluators to obtain our test set. Since answers are subjective and noisy, we labeled them twice, making sure the same evaluator does not receive the same question twice. After that, we manually reviewed the answers for semantic disagreements and mistakes. We relabeled such samples two more times and fixed the disagreements, preferring the semantics of the majority of responders but preserving maximal variety in the responses. Finally, we condensed this into 1k high-quality answers to 500 questions, with two correct but diverse answers per question. The dataset evaluates a range of competencies, including action and justification, attention, description, localisation, identification, counting and anticipation, as shown in Figure 2.
@@ 3.2 Evaluation Metric
Table 1: Lingo-Judge Performance. Correlation with human ratings, validation accuracy, and time taken to run of our proposed LingoQA evaluation metric compared to previous language-based metrics. All metrics use textual ground truth and have no access to vision information. Further examples are presented in Appendix B.
GPT-4 based evaluation
Inspired by the G-Eval metric [36], we used GPT-4 to evaluate answers on a larger scale. Given a question and answer pair from the test set and a model’s answer, we ask GPT-4 to evaluate whether the model’s answer corresponds to a human’s answer. Notice that it does not make use of any visual input. We experimented with prompts and methods achieving good quality of judgements. We achieved the highest accuracy by employing chain-of-thought prompting where we ask GPT-4 to first come up with an evaluation strategy before grading a model’s answer. However, as shown in Table 1, this leads to increased inference time. Further details are provided in Appendix C. Unfortunately, we found GPT-4 based evaluation impractical to use as a main development and training metric due to the time required to evaluate answers on our relatively small evaluation dataset (from 13min up to 50min for a single evaluation due to the API rate limit).
Lingo-Judge
@@
In comparison to metrics such as CIDEr, which provide a system-level performance metric, the classifier provides a probability of correctness for each of the model predictions, meaning that it provides metrics at the sample level. Examples are provided in the Appendix B. This means that 100% classifier accuracy is easy to interpret. The classifier allows us to compute metrics during training, running over our full evaluation dataset in 10 seconds using an A100 GPU.
Correlation to human ratings
We studied empirical correlation of various metrics with human judgments. Several human annotators assigned a scalar score [0, 1] to the inference outputs of 17 different models which can be interpreted as the probability that the response correctly addresses the question [34]. Notably, this process takes several days, highlighting the need for an automated evaluation metric that provides faster development feedback. The final human score of each model is the average of all inference output scores. Further details regarding the methodology for the correlation analysis are in the Appendix D.
@@
Scenarios QA pairs QA per scenario
Action 24.5k 267.8k $\approx 10.9$
Scenery 3.5k 152.5k $\approx 43.6$
Eval. Dataset 100 1000 $10$
Table 2: Dataset Split. It consists of three different datasets of varying annotation densities. The Action dataset focuses on questions related to driving behaviours, the Scenery dataset focuses on perception capabilities, while the evaluation dataset is designed to probe a range of competencies.
@@
Table 3: Dataset Features. The dataset that we introduce alongside our benchmark consists of questions related to object presence, as well as action, justification, attention, localisation, counting, anticipation, and counterfactuals. In total, it has a similar size to other driving-related datasets such as NuScenesQA, while having a much higher diversity and not being limited to questions related to object positioning.
3.3 Datasets
Figure 2: Dataset Statistics. Dataset split by the number of question-answer pairs for the competencies covered and for the objects referred. One question-answer pair might cover more than one competency or object, hence the total is higher than the size of the datasets. The Action and Scenery datasets have complementary strengths, with one focused more on action-justification competencies and one more on description and localisation.
We created a collection of datasets for bringing language to autonomous driving. The total dataset size is 419.9k question-answer pairs, where a single data sample consists of a 4-second video clip at 1Hz. The total size of the dataset is about 10x larger than BDD-X [29], as shown in Table 3. Compared to prior datasets such as NuScenesQA [42], our dataset contains reasoning pairs in addition to object presence, description, and localisation. The answers are also more free-form and more complex, with an average answer length of 17.2 words versus 1.0 words in NuScenesQA. Examples of question answers pairs from LingoQA are shown in Appendix A.
Our labeled autonomous driving training dataset consists of two complementary parts: the action dataset and the scenery dataset.
@@ 3.3 Datasets
This dataset was made from a recorded driving corpus of interesting events where the car’s behavior changes, such as decelerations, accelerations, lane changes, narrow gaps, and turns. Such events were succinctly labeled by driving operators with very short high-level descriptions of the situations and behavioral policies (e.g. “following lane, pedestrian on a zebra crossing, should stop”). Additionally, we added metadata for such events from various perception systems, such as traffic light presence, vehicles and pedestrian visual detectors, weather descriptors, as well as other metadata (speed, steering wheel position, and road type from the map data). Using this data, we developed prompt templates for (1) describing the current action and its justification and (2) a set of example questions and hints about what the answer should mention. Next, we used those prompts with GPT-3.5 to rephrase, answer, and extend the example questions using the provided action description and answer hints. We rebalanced events by bucketing by actions and behavioral policies and sampled up to 500 events from each bucket without replacement, leading to 24,577 video snippets with 167,774 question/answer pairs.
Scenery dataset
The scenery dataset was built to complement the action dataset by focusing on perception-related questions in addition to driving behaviours. The dataset was made by densely and thoroughly labelling three 30-minute driving sessions with the ELAN video annotation software[38]. For the entire duration of the driving sessions, we provided short captions in about 15 different categories:
  • Driver’s actions, and their justifications
  • Driver’s attention
  • Observations about relevant vehicles, pedestrians, and other road actors with their visual descriptions
  • Observations about relevant road elements such as traffic lights, traffic islands, lane and intersection structures
  • Miscellaneous observations about the environment, such as weather, tube stations, and buildings.
Then, for every keyframe every second (1fps), we collect all annotations around this frame and build a textual description containing the driver’s actions and their justifications, the objects requiring the driver’s attention, and the observations. As opposed to the Action dataset, where recommended questions were provided to GPT-3.5 for rephasing, for the Scenery dataset, we asked GPT-4 to generate questions and answers using a set of generic prompts, but also using a prompt with the chain of thought specifically targeting perception questions. This forced GPT-4 to generate many diverse questions and answers. This led to a high-quality diverse dataset with about 43 QA-pairs per video.
@@ 4.1 Architecture
We leverage pretrained LLMs to give LingoQA Baseline the ability to answer general questions related to both driving scenes, as well as general knowledge. We use Vicuna v1.5 [13] with 7B parameters built on top of Llama-2 [47]. The language model is auto-regressive and hence can be conditioned on textual inputs, as well as image tokens. The training objective is to predict the next language token in a sequence. We mask all tokens from the training loss that belong to the text prompt, including question and chat history.
@@ 4.2 Training Recipe
Our training uses a two-step approach to better utilise video features and improve learning when answering questions based on video data. Through this two-step training, we aim to better understand and use video data. In the first stage, we train the self-attention layers for the LLM and the vision encoder (QKV), together with all the Q-Former parameters and the linear language projection layer. In the second stage, we fine-tune the same parameters as in the previous stage, keeping the vision encoder frozen. Further details regarding training parameters are presented in Appendix F.
@@
Stage 1: Pre-training for feature alignment.
In the first stage, we pre-train the model on the GQA and SVIT datasets to align image features with the embedding space of the pretrained LLM. The GQA dataset [23] contains more than 22M questions over 113k images. The recently introduced SVIT dataset [61] contains 4.2M question-answer pairs over 108.1k images. We leverage initial weights from different models to accelerate the training process. We initialise the vision encoder using publicly available weights of OpenCLIP [24], the Q-Former from BLIP2 weights [31], and language model from Vicuna v1.5 [13].
Stage 2: Fine-tuning for video QA.
In the second stage, we fine-tune the model on our video question-answering Action and Scenery datasets described in Section 3.3. During the fine-tuning phase, each sample is composed of 5 frames taken from a 4-second span of video, accompanied by a QA-pair. To facilitate further exploration of autonomous driving QA, we open-source the dataset used to fine-tune LingoQA Baseline.
5 Empirical Evaluation on LingoQA

### v4

# https://wayve.ai/science/

@@ Abstract
We introduce LingoQA, a novel dataset and benchmark for visual question answering in autonomous driving. The dataset contains 28K unique short video scenarios, and 419K annotations. Evaluating state-of-the-art vision-language models on our benchmark shows that their performance is below human capabilities, with GPT-4V responding truthfully to 59.6% of the questions compared to 96.6% for humans. For evaluation, we propose a truthfulness classifier, called Lingo-Judge, that achieves a 0.95 Spearman correlation coefficient to human evaluations, surpassing existing techniques like METEOR, BLEU, CIDEr, and GPT-4. We establish a baseline vision-language model and run extensive ablation studies to understand its performance. We release our dataset and benchmark [^1] https://github.com/wayveai/LingoQA as an evaluation platform for vision-language models in autonomous driving.
Keywords:
Autonomous Driving Question Answering Vision-language
*Equal contributions
Figure 1: Examples from the LingoQA benchmark designed to evaluate whether models can describe the elements in the scene (description), predict what future action should be taken (action) and why (justification) and explain what they are paying attention to (attention).
@@ 1 Introduction
Our focus is on vision-only end-to-end autonomous driving, aiming to bridge the gap between data-driven decision-making and user trust. We introduce LingoQA, a benchmark designed for autonomous driving VQA, utilizing a novel dataset comprising more than 419k QA pairs. Distinguished by its free-form approach to questions and answers, this dataset broadens the scope of autonomous driving QA, encompassing reasoning and action justifications. Additionally, we publish a comprehensive evaluation suite consisting of 1,000 examples. At the core of our benchmark lies a novel evaluation metric based on a learned text classifier called Lingo-Judge, inspired by GPT-Judge used in TruthfulQA [34]. We perform rigorous studies correlating automatic metrics to human preferences and find that Lingo-Judge achieves a 0.950 Spearman and 0.993 Pearson correlation coefficient, surpassing existing automated labelling techniques like METEOR [5], BLEU [39], CIDEr [49], and GPT-4 [38] on our benchmark, while being fast enough for frequent runs during training and development. The evaluation code and the weights for the classifier will be released with the paper to support robust benchmarking of visual question answering in autonomous driving.
@@
  • LingoQA Benchmark: We introduce a novel benchmark for autonomous driving VQA using a learned text classifier for evaluation. It outperforms existing metrics, including GPT-4, with a Spearman coefficient of 0.950 indicating a strong correlation with human evaluation.
  • LingoQA Dataset: Our 419.9k QA pair dataset stands out with its free-form questions and answers, covering not just perception but also driving reasoning from the drivers directly, broadening the scope of autonomous driving datasets.
  • LingoQA Baseline: Through testing of various vision-language components on LingoQA, we find that the most effective approach involves partially fine-tuning the attention layers of our vision-language model equipped with Vicuna-1.5-7B [13] and a late video fusion technique. We establish a new baseline for this field with an identified model combination. Example outputs from the model are shown in Figure 1.
@@ 2.2 Evaluation Metrics
Progress has been relatively slow in developing vision-language models for autonomous driving, with only a few works aiming to quantitatively improve upon prior work [28, 27, 56]. A key challenge consists of automated, reproducible evaluation metrics that are highly correlated with human ratings, particularly due to the inherent complexities in assessing natural language. ADAPT [27] reports human feedback in addition to standard natural language generation metrics, while DriveGPT [56] reports ChatGPT ratings. Automated methods such as BLEU [39], METEOR [5], ROUGE [33] show weak alignment with human feedback [49]. CIDEr [49] is also based on n-gram level similarity, as opposed to capturing the correctness of an answer based on its meaning. Newer evaluation metrics using ChatGPT have shown improvement in the area of sentence understanding, while still having limitations, such as providing high scores to elaborate, eloquent, but incorrect sentences [2]. Evaluation based on human feedback is subjective and difficult to reproduce. In this work, we address this challenge by introducing a novel VQA benchmark for autonomous driving that checks for factual correctness and is highly correlated to human correctness ratings on our proposed evaluation dataset.
@@ 2.3 Datasets for Autonomous Driving
Our proposed dataset LingoQA addresses the existing gap in autonomous driving as it contains a diverse set of questions related to driving behaviours and scenery in addition to perception questions related to object presence and positioning. The evaluation dataset probes areas such as description, counting, localisation, anticipation, attention, and action justification. This dataset has the strength of being diverse with respect to the language used while being grounded in human reasoning. Examples of the complex questions and answers existent in the dataset are provided in Appendix 0.A.
@@ 3 LingoQA Benchmark
In this section, we introduce LingoQA, a benchmark to evaluate autonomous driving reasoning through video question-answering. The benchmark consists of an automated evaluation metric and the corresponding datasets for evaluation and fine-tuning.
@@ 3.1 Evaluation Metric
N-Gram Matching Metrics.
Common language-based metrics such as BLEU [39], METEOR [5] and CIDEr [49] are still the most common metrics used for evaluating captioning in autonomous driving, despite their known limitations, such as relying heavily on the n-gram frequency as opposed to the underlying meaning of the answer. The BLEU [39] metric computes the n-gram based precision of the candidate sequence compared to the reference sequence. Similarly, ROUGE [33] computes the recall-based n-gram, while METEOR [5] computes the F-measure. CIDEr [49] is based on using TF-IDF to provide a lower weight to terms that are commonly reported in the corpus.
Table 1: Lingo-Judge performance. Correlation with human ratings, validation accuracy, and time taken to run of our proposed LingoQA evaluation metric compared to previous language-based metrics. All metrics use textual ground truth and have no access to vision information. Further examples are presented in Appendix 0.B.
@@
GPT-4 based evaluation.
Inspired by the G-Eval metric [36], we used GPT-4 to evaluate answers on a larger scale. Given a question and answer pair from the test set and a model’s answer, we ask GPT-4 to evaluate whether the model’s answer corresponds to a human’s answer. Notice that it does not make use of any visual input. We experimented with prompts and methods achieving good quality of judgements. We achieved the highest accuracy by employing chain-of-thought prompting where we ask GPT-4 to first come up with an evaluation strategy before grading a model’s answer. However, as shown in Table 1, this leads to increased inference time. Further details are provided in Appendix 0.C. Unfortunately, we found GPT-4 based evaluation impractical to use as a main development and training metric due to the time required to evaluate answers on our relatively small evaluation dataset (from 13min up to 50min for a single evaluation due to the API rate limit).
Lingo-Judge.
@@
In comparison to metrics such as CIDEr, which provide a system-level performance metric, the classifier provides a probability of correctness for each of the model predictions, meaning that it provides metrics at the sample level. Examples are provided in the Appendix 0.B. This means that 100% classifier accuracy is easy to interpret. The classifier allows us to compute metrics during training, running over our full evaluation dataset in 10 seconds using an A100 GPU.
Correlation to human ratings.
We studied empirical correlation of various metrics with human judgments. Several human annotators assigned a scalar score [0, 1] to the inference outputs of 17 different models which can be interpreted as the probability that the response correctly addresses the question [34]. Notably, this process takes several days, highlighting the need for an automated evaluation metric that provides faster development feedback. The final human score of each model is the average of all inference output scores. Further details regarding the methodology for the correlation analysis are in the Appendix 0.D.
@@
3.2 Datasets
We created a collection of datasets for bringing language to autonomous driving. The total dataset size is 419.9k question-answer pairs, where a single data sample consists of a 4-second video clip at 1Hz. The total size of the dataset is about 10x larger than BDD-X [29], as shown in Table 2. Compared to prior datasets such as NuScenesQA [42], our dataset contains reasoning pairs in addition to object presence, description, and localisation. The answers are also free-form and more complex, with an average answer length of 17.2 words versus 1.0 words in NuScenesQA. Examples of question answers pairs from LingoQA are shown in Appendix 0.A.
Table 2: Dataset comparison. The dataset that we introduce has a similar size to other driving-related while having a much higher diversity. A scenario refers to a unique short video sequence, that may have multiple annotations attached to it.
@@ 3.2 Datasets
Our labeled autonomous driving training dataset consists of two complementary parts: the action dataset and the scenery dataset. The breakdown of the datasets is shown in Table 3.
@@
The action dataset was created from recorded driving sessions featuring notable events where the car’s behavior changes, such as decelerations, accelerations, lane changes, narrow gaps, and turns. These events were labeled by driving operators with concise, high-level descriptions of the situations and behaviors (e.g., "following lane, pedestrian on a zebra crossing, should stop"). Additional metadata was included from various perception systems, covering traffic light presence, vehicle and pedestrian detection, weather conditions, and other data like speed, steering wheel position, and road type. Using this information, we developed prompt templates for (1) describing the current action and its justification and (2) generating example questions and hints for expected answers. We then used these prompts with GPT-3.5 to rephrase, answer, and extend the example questions based on the provided action descriptions and answer hints. Events were rebalanced by categorizing them based on actions and behavioral policies, and up to 500 events were sampled from each category, resulting in 24,577 video snippets with 167,774 question/answer pairs.
Scenery dataset.
The scenery dataset was designed to complement the action dataset by focusing on fine-grained perception-related questions. It was created by densely and thoroughly annotating three 30-minute driving sessions using the ELAN video annotation software. [40] During these sessions, brief captions were provided in approximately 15 different categories, including driver actions, justifications, attention, observations of vehicles, pedestrians, road elements (such as traffic lights, traffic islands and intersections), and environmental details (like weather and buildings). Annotations were collected every second (1 fps) for every frame to build textual descriptions that included driver actions, justifications, and observations. Unlike the Action dataset, where GPT-3.5 was used only for question rephrasing, GPT-4 was employed for the Scenery dataset to generate questions and answers. GPT-4 was prompted with chain of thought specifically targeting perception questions, resulting in a diverse and high-quality dataset with about 43 QA pairs per video.
Dataset statistics.
@@
Table 3: Dataset split. The Action dataset focuses on questions related to driving behaviours, the Scenery dataset focuses on perception capabilities, while the evaluation dataset is designed to probe a range of competencies.
Scenarios QA pairs QA per scenario
Action 24.5k 267.8k $\approx 10.9$
Scenery 3.5k 152.5k $\approx 43.6$
Eval. Dataset 100 1000 $10$
Figure 2: Fine-tuning dataset statistics. The Action and Scenery datasets have complementary strengths, one focused on action-justification competencies and one on description and localisation. The number of question-answer (QA) pairs for the competencies covered and for the objects referred are shown. One QA pair might cover more than one competency or object, hence the total is higher than dataset size.
Evaluation dataset.
For evaluation we collected a small, low-density dataset from in-house human labelers, creating both the questions and the answers associated with the short videos. We labeled a small portion of held-out data on 500 human-generated questions using 20+ different evaluators to obtain our test set. Since answers are subjective and noisy, we labeled them twice, making sure the same evaluator does not receive the same question twice. After that, we manually reviewed the answers for semantic disagreements and mistakes. We relabeled such samples two more times and fixed the disagreements, preferring the semantics of the majority of responders but preserving maximal variety in the responses. Finally, we condensed this into 1k high-quality answers to 500 questions, with two correct but diverse answers per question, as shown in Table 3. The dataset evaluates a range of competencies, including action and justification, attention, description, localisation, identification, counting and anticipation, as shown in Appendix 0.A.
@@ 4.1 Architecture
We leverage pretrained LLMs to give LingoQA Baseline the ability to answer general questions related to both driving scenes, as well as general knowledge. We use Vicuna v1.5 [13] with 7B parameters built on top of Llama-2 [47]. The language model is autoregressive and hence can be conditioned on textual inputs, as well as image tokens. The training objective is to predict the next language token in a sequence. We mask all tokens from the training loss that belong to the text prompt, including question and chat history.
@@ 4.2 Training Recipe
Our training uses a two-step approach to better utilise video features and improve learning when answering questions based on video data. Through this two-step training, we aim to better understand and use video data. In the first stage, we train the self-attention layers for the LLM and the vision encoder (QKV), together with all the Q-Former parameters and the linear language projection layer. In the second stage, we fine-tune the same parameters as in the previous stage, keeping the vision encoder frozen. Further details regarding training parameters are presented in Appendix 0.F.
Stage 1: Pre-training for feature alignment.
In the first stage, we pre-train the self-attention layers of the LLM and the vision encoder (QKV) on the GQA and SVIT datasets to align image features with the embedding space of the pretrained LLM. The GQA dataset [23] contains more than 22M questions over 113k images. The recently introduced SVIT dataset [61] contains 4.2M question-answer pairs over 108.1k images. We leverage initial weights from different models to accelerate the training process. We initialise the vision encoder using publicly available weights of OpenCLIP [24], the Q-Former from BLIP2 weights [32], and language model from Vicuna v1.5 [13].
Stage 2: Fine-tuning for video QA.
In the second stage, we fine-tune the model on our video question-answering Action and Scenery datasets described in Section 3.2. During the fine-tuning phase, each sample is composed of 5 frames taken from a 4-second span of video, accompanied by a QA-pair. The dataset used to fine-tune LingoQA Baseline is open-sourced.
5 Empirical Evaluation on LingoQA
@@
5.1 Ablation Studies on LingoQA